Sarah Meeham blends coffee for tasti delight. she needs to prepare 140 pounds of blended coffee beans selling for $4.50 per pund. She plans to do thid by blending together a hig-quality bean costing $5.50 per pound and a chaper bean at $3.50 per pound. To the nearest pound, find how much high-quality coffee bean and how much cheaper coffee bean she should blend
Answer by richwmiller(17219)

a+b=140,
5.5*a+3.5*b=4.5*140
a=140-b
5.5*(140-b)+3.5*b=630
770-5.5b+3.5*b=630
-2*b=-140
b=70
a=140-b
a=70 b=70
check
5.5*70+3.5*70=4.5*140
385+245=630
630=630
ok